OpenStudy (anonymous):

Given the equation 5x + y = 7, which equation below would cause a consistent-independent system? 10x + 2y = 14 -15x - 3y = -6 5x + y = -7 6x + 2y = 7

OpenStudy (anonymous):

help! please!

jimthompson5910 (jim_thompson5910):

5x + y = 7 and 5x + y = -7 will produce an inconsistent system this is because you'll have 2 parallel lines when you graph them leading to no intersections (no intersections ---> no solutions)

jimthompson5910 (jim_thompson5910):

so choice C is out

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

10x + 2y = 14 2(5x + y) = 14 5x + y = 14/2 5x+y = 7 So choice A is equivalent to the given equation. Therefore, we'll get a consistent dependent system. Choice A is out.

jimthompson5910 (jim_thompson5910):

I'll let you do check the others

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

is it choice D?

jimthompson5910 (jim_thompson5910):

so you eliminated choice B? what did you get for B?

OpenStudy (anonymous):

for choice B was parallel to 5x + y = 7

jimthompson5910 (jim_thompson5910):

correct, so B is out D is the only thing left and it is indeed the answer

OpenStudy (anonymous):

ok thank you very much :D

jimthompson5910 (jim_thompson5910):

you're welcome